OpenStudy (anonymous):

medal! & i'll fan everyone who answers. one number is four more than another, and their sum is 60. what is the smaller number? A. 26 B. 28 C. 32

OpenStudy (anonymous):

wanna do algebra or make a guess?

OpenStudy (anonymous):

or just check the possible answers to see which one is right? we could do that too, probably easiest

OpenStudy (anonymous):

I don't know how to do it

OpenStudy (anonymous):

what is 4 more than 26?

OpenStudy (anonymous):

i.e. what is \(26+4\)?

OpenStudy (anonymous):

30

OpenStudy (anonymous):

ok good now what is \(24+30\) ?

OpenStudy (anonymous):

54. so the correct choice is B?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Smaller number is 28.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

now you know how to do it

OpenStudy (anonymous):

60 = s + (s+4) 60 = 2s + 4 56 = 2s smaller number = 28

OpenStudy (anonymous):

OK sorry, that might be confusing.